# TIF-UNIMI-2025-23

### v1

@@ 1 Soft resummation from QED to QCD
The all-order resummation of logarithmically enhanced contributions due to the emission of gauge vector bosons, and specifically its use in the context of QCD, is both a standard topic and an active research field. The underpinnings of these techniques in classic QED results are covered both in old [Berestetskii:1982qgu] and more recent [Sterman:1993hfp] graduate-level textbooks. Some recent developments, such as for example webs [White:2015wha] have been reviewed, and the approach to resummation based on soft-collinear effective theories has been the subject of textbook treatments [Becher:2014oda]. However, a treatment at the introductory level of the direct QCD approach, rooted in the original papers of Sterman [Sterman:1986aj], Catani and Trentadue [Catani:1989ne] is rather more difficult to find. It is the purpose of these elementary lectures to provide such an introduction.
In order to make our treatment accessible and self-contained, we start by reviewing textbook [Peskin:1995ev] material on the factorization of soft and collinear emission, discussing the cancellation of infrared singularities and the factorization of collinear singularities. We then proceed to show the origin of double-logarithmically enhanced contributions, and their all-order exponentiation. We take particular care in explaining some simple computational tricks that are used in order to obtain the desired results, for instance in the treatment of phase space. We then review basic ideas on renormalization group invariance, the Callan-Symanzik renormalization group equation and its solution, again at a basic textbook level. We then show how the standard threshold resummation [Sterman:1986aj, Catani:1989ne] can be derived using a renormalization-group argument [Contopanagos:1996nh, Forte:2002ni], explain the underlying physical argument, illustrate the predictive power of the resummed results, and demonstrate the equivalence of various forms that it can take. We conclude by briefly touching upon transverse momentum resummation.
@@
Soft resummation in QCD sums to all perturbative orders contributions that are enhanced by powers of the logarithm of the ratio of a soft scale and the hard scale of the process. They arise because contributions coming from real gluon emission diverge in the threshold limit in which the soft scale vanishes. The divergence is either at threshold by virtual corrections, in which the emitted gluon is absent, and the log is the left-over of the cancellation away from threshold. This log gets combined with a second log arising from the region of integration over the transverse momentum of the emitted gluon in which the emitted and emitting parton momenta become collinear (i.e. parallel). Resummation is possible because of the factorized nature of the gluon emission process when the gluon is soft (i.e. its energy tends to zero) or collinear (i.e. its emission angle with respect to the emitting parton tends to zero). Multiple emissions can thus be treated as a branching process with a suitable factorization of phase space. We discuss in turn the factorized nature of soft and collinear emission; singularities in the soft limit and the logs they leave behind; and the exponentiation of multiple emission.
@@ 2 Factorized emission
The emission of soft gluons and the emission of collinear partons (quark or gluons) from either quarks or gluons factorizes, in the sense that the process with an extra gluon or parton in the final state can be written in terms of that without it, times a universal emission kernel. The factorization happens at the level of amplitudes for soft gluons, and at the level of squared amplitudes for collinear partons — factorization of collinear emission at the amplitude level is rather more subtle and goes beyond the scope of our discussion (see Ref. [Duhr:2025cye]). We summarize main results and refer to textbooks [Peskin:1995ev, Ellis:1996mzs] and reference works [Collins:2011zzd, Weinberg:1995mt] for a more detailed discussion.
@@ 2.1 Collinear emission
We assume for definiteness that the gluon is radiated by the incoming quark, and we follow notation and conventions of Ref. [Peskin:1995ev], to which we also refer for the Feynman rules of QCD. We can write the corresponding contribution to the amplitude $M_{1}$ for the process with the extra gluon in terms of the tree-level amplitude $M_{0}$ as
@@
An analogous calculation yields the contribution of the emission from the antiquark line, $M_{1}^{\bar{q}}$. We now note that in the collinear limit the momentum in the intermediate propagator goes on shell. Indeed, we have
@@
where it is understood that the argument $p_{1}-k$ of the spinor $u_{s}$ is an on-shell momentum, $(p_{1}-k)^{2}=0$.
@@
Now, it should be clear that nowhere in the above argument the explicit form of $M_{0}$ was used. So what we have actually proven is that radiation of a gluon from an external quark line produces a logarithmically enhanced contribution to the square amplitude which is proportional to the original amplitude, but with an incoming quark momentum rescaled by a factor $z\leq 1$, and a universal splitting function which only depends on the radiation process. A similar argument also works for gluon radiation from an incoming gluon line, because one may likewise rewrite in the collinear limit the numerator of the intermediate gluon propagator as a sum over on-shell gluon polarization vectors, up to subtleties related to the choice of gauge in the nonabelian case. Of course, in this case the splitting function will be a different $P^{r}_{gg}$ now depending on the $g(p_{1})\to g(p_{1}-k)g(k)$ matrix element. Furthermore, what is being used in the argument is the on-shell behavior of the intermediate propagator. This means that the argument also works in the off-diagonal case, except that in this case an incoming quark is replaced by a gluon with factor $P^{r}_{gq}(z)$ and an incoming gluon is replaced by a quark with factor $P^{r}_{qg}(z)$. Finally, it is clear that the argument also works for massive quarks, though in this case the logarithmic integral at the lower end is regulated by the quark mass.
@@ 3.1 The Sudakov double log
In order to understand the origin of the double soft-collinear log, consider again the full set of gluon emission corrections to the tree-level Drell-Yan process (see Fig. 3). We are specifically interested in the form of the phase space integration over the emitted gluon momentum $k$. To this purpose, it is convenient to use a slightly different version of the Sudakov parametrization, namely the light-cone parametrization, in which instead of $p_{1}$, and $p_{1}$ we adopt as basis vectors their linear combinations
@@
Because we are interested in the limit in which $k$ is both collinear and soft, we can use the eikonal approximation Eq. (17). We include all first-order real emission corrections, i.e. those corresponding both to radiation from the quark and the antiquark, because the inclusion of all gauge-invariant set of diagrams avoids some subtleties that would instead arise when computing directly the splitting function corresponding to emission from a single quark or antiquark line as in Sect. 2.1-2.2. Of course, the full set of diagrams also involves virtual loop corrections that to this order interfere with the leading-order process: we will come to these in Sect. 3.3
@@
$M_{1}=M_{1}^{q}+M_{1}^{\bar{q}}=M_{0}t^{a}g_{s}\left(\frac{p_{1}^{\mu}}{p_{1}\cdot k}-\frac{p_{2}^{\mu}}{p_{2}\cdot k}\right)\epsilon^{*}_{\mu}(k)\left[1+O(1-x)\right],$ (25)
@@
$|M_{1}|^{2}=|M_{0}|^{2}C_{F}g^{2}_{s}\frac{2p_{1}\cdot p_{2}}{p_{1}\cdot k\,p_{2}\cdot k}\left[1+O(1-x)\right],$ (26)
@@
Using Eq. (32) it is apparent that only the second term in Eq. (30) contributes to the integral, which is then straightforwardly giving a square (usually called double) log. The upper limit of integration is the maximum allowed value of $k_{t}^{2}$, which is proportional to the center-of-mass energy $s$, while at the lower limit the integral diverges, so we must introduce some infrared cutoff $\mu^{2}$ in order to make sense of it. The treatment of this double divergence - one collinear and one soft - will be the subject of the next two sections. We thus get
$\int d\Phi_{k}|M_{1}|^{2}=C_{F}|M_{0}|^{2}\frac{\alpha}{2\pi}\ln^{2}\frac{s}{\mu^{2}}+\hbox{less singular}$ (33)
where we have neglected all contributions that are less singular as $\mu^{2}\to 0$. This includes both the O(1x)O(1-x) corrections to the eikonal approximation, which would in fact be suppressed by inverse powers of the dimensionful variable $s$ which contains the large scale of the process, but also lower powers of log, which arise when expressing the maximum value of $k_{t}^{2}$ in terms if $s$. For example, for the simple $2\to 2$ process of Fig. 1 one has ${k_{t}^{2}}_{\rm max}=s/4$ so $\ln^{2}{k_{t}^{2}}_{\rm max}=\ln^{2}s-\ln 4\ln s+\text{non log}$.
@@ 3.2 Mass singularities and their factorization
The same conclusion holds if the radiating parton is off-shell, because then the denominator of the intermediate propagator Eq. (10) contains a contribution equal to the off-shellness $p_{1}^{2}$, and thus it does not vanish when $k_{t}^{2}\to 0$.
@@ 3.3 Cancellation of infrared singularities
After subtraction of the collinear singularity - which has been factorized into the PDF, where it is then regulated by the nonperturbative physics which determines the PDF - we are still left with an infrared singularity. This can be seen clearly by looking at the explicit form of the splitting function $P^{r}_{qq}$:
@@
A classic result of quantum electrodynamics, the Kinoshita-Lee-Nauenberg (KLN) theorem, guarantees that once one consistently includes loop corrections along with real emission contributions this singularity cancels. The argument generalizes in QCD for colorless final states [Weinberg:1995mt]. The cancellation is illustrated graphically in Fig. 4. The argument of Sect. 3.1 shows that the infrared divergent contribution to the cross-section can be obtained from the interference of the contributions to the amplitude coming from gluons emitted from different quark lines lines. Indeed, the infrared singularity comes from the $x\to 1$ limit of the denominator in the amplitude Eq. (28), which in turn arises from the interference of the two eikonal factors proportional to $p_{1}^{\mu}$ and $p_{2}^{\mu}$ in Eq. (25). The KLN theorem ensures that this singularity cancels against an equal and opposite singularity which is present in the virtual contribution that corresponds to the gluon loop obtained by connecting the endpoints of the two interfering emitted real gluons. This loop correction has of course the same final state particles as the amplitude without either of the two interfering final-state real emitted gluons, and thus it interferes with it when calculating the square modulus of the amplitude, thereby producing a contribution of the same order of the square of either of the real emission terms (see Figure).
@@
Hence the partonic cross-section contains a double-logarithmic contribution, coming from interplay between the leftover of the infrared singularity, that produces a $\frac{1}{1-z}$ behavior, and the leftover of the collinear singularity, that produces a $\ln{k_{t}^{2}}_{\rm max}$ behavior, which in turns is sensitive to the infrared scale $Q^{2}(1-z)^{2}$.
@@ 4.1 Momentum conservation and phase space factorization
In a multiple emission process an incoming parton with momentum $p_{1}$ emits a first gluon with momentum $k_{1}$, the outgoing parton after the first emission, with momentum given by $p_{1}-k$ is then the incoming parton for the second emission, and so forth. Using the Sudakov parametrization Eq. (2) it is clear that after the first emission of a gluon with momentum
@@ 5 Renormalization group basics
The basic idea of renormalization is that in a quantum field theory all couplings and fields must be defined in terms of a reference scale, in such a way that all quantum fluctuations at distances shorter than that of the reference scale are absorbed in the definition of the couplings and fields. The fields and couplings must consequently be defined by relating them to a reference scale. The quantum field theory then provides relations between observable quantities, with the couplings and fields used as a way to relate observables to each others. Infinities only appear if one tries to express observable quantities in terms of unobservable parameters defined at infinitely short distance scales.
For example, one may define the charge of the electron in quantum electrodynamics in terms of the Coulomb force at large distance, i.e. at the scale of the (rest) electron mass. The (say) elastic electron-electron scattering cross-section is finite once re-expressed in terms of the electron charge thus defined. Renormalization group invariance is then the consequence of the fact that physical predictions do not depend on the renormalization scale that has been chosen to define the couplings. Because the dependence of the couplings on this scale is a universal property of the theory, this independence expresses the scale dependence of physical observables in terms of this universal scale dependence of the coupling.
The basic idea of deriving collinear exponentiation from renormalization group invariance, that we will present shortly, takes this one step further, based on the observation that in QCD physically observable cross-sections are expressed in factorized form in terms of parton distributions, convoluted with partonic cross-sections. This factorization requires introducing a scale, as discussed in Sect. 3.2 for the subtraction of collinear singularities. As discussed in Sect. 4.2, this can be viewed as a scale that splits in two the parton radiation process, with radiation below the scale included in the parton distribution, and radiation above it included in the hard matrix element. Physical predictions are independent of this scale, hence this scale dependence must cancel between the PDF and the hard cross-section: indeed we have seen that it is expressed in terms of a universal, process-independent splitting function, or rather, its Mellin transform, the anomalous dimension Eq. (45). But the PDF and hard cross-section depend on different variables: the hard cross-section depends on the specific hard scale of the process, while the PDF does not. As we shall see, dimensional analysis then implies that the cancellation of the dependence on the factorization scale can only go through the scale dependence of the coupling constant: it is calculable in terms of it, and the universal anomalous dimension. It can then be integrated away, leading to all-order resummation and exponentiation of logs of the hard physical scale.

### v2

@@ 1 Soft resummation from QED to QCD
The all-order resummation of logarithmically enhanced contributions due to the emission of gauge vector bosons, and specifically its use in the context of QCD, is both a standard topic and an active research field. The underpinnings of these techniques in classic QED results are covered both in old [2] and more recent [16] graduate-level textbooks. Some recent developments, such as for example webs [18] have been reviewed, and the approach to resummation based on soft-collinear effective theories has been the subject of textbook treatments [1]. However, a treatment at the introductory level of the direct QCD approach, rooted in the original papers of Sterman [15], Catani and Trentadue [4] is rather more difficult to find. It is the purpose of these elementary lectures to provide such an introduction.
In order to make our treatment accessible and self-contained, we start by reviewing textbook [14] material on the factorization of soft and collinear emission, discussing the cancellation of infrared singularities and the factorization of collinear singularities. We then proceed to show the origin of double-logarithmically enhanced contributions, and their all-order exponentiation. We take particular care in explaining some simple computational tricks that are used in order to obtain the desired results, for instance in the treatment of phase space. We then review basic ideas on renormalization group invariance, the Callan-Symanzik renormalization group equation and its solution, again at a basic textbook level. We then show how the standard threshold resummation [15, 4] can be derived using a renormalization-group argument [6, 9], explain the underlying physical argument, illustrate the predictive power of the resummed results, and demonstrate the equivalence of various forms that it can take. We conclude by briefly touching upon transverse momentum resummation.
@@
Soft resummation in QCD sums to all perturbative orders contributions that are enhanced by powers of the logarithm of the ratio of a soft scale and the hard scale of the process. They arise because contributions coming from real gluon emission diverge in the threshold limit in which the soft scale vanishes. The divergence is canceled at threshold by virtual corrections, in which the emitted gluon is absent, and the log is the left-over of the cancellation away from threshold. This log gets combined with a second log arising from the region of integration over the transverse momentum of the emitted gluon in which the emitted and emitting parton momenta become collinear (i.e. parallel). Resummation is possible because of the factorized nature of the gluon emission process when the gluon is soft (i.e. its energy tends to zero) or collinear (i.e. its emission angle with respect to the emitting parton tends to zero). Multiple emissions can thus be treated as a branching process with a suitable factorization of phase space. We discuss in turn the factorized nature of soft and collinear emission; singularities in the soft limit and the logs they leave behind; and the exponentiation of multiple emission.
@@ 2 Factorized emission
The emission of soft gluons and the emission of collinear partons (quark or gluons) from either quarks or gluons factorizes, in the sense that the process with an extra gluon or parton in the final state can be written in terms of that without it, times a universal emission kernel. The factorization happens at the level of amplitudes for soft gluons, and at the level of squared amplitudes for collinear partons — factorization of collinear emission at the amplitude level is rather more subtle and goes beyond the scope of our discussion (see Ref. [7]). We summarize main results and refer to textbooks [14, 8] and reference works [5, 17] for a more detailed discussion.
@@ 2.1 Collinear emission
We assume for definiteness that the gluon is radiated by the incoming quark, and we follow notation and conventions of Ref. [14], to which we also refer for the Feynman rules of QCD. We can write the corresponding contribution to the amplitude $M_{1}$ for the process with the extra gluon in terms of the tree-level amplitude $M_{0}$ as
@@
where $t^{a}$ are SU(3) gauge group generators, and $\epsilon^{a}_{\mu}$ is the gluon polarization. An analogous calculation yields the contribution of the emission from the antiquark line, $M_{1}^{\bar{q}}$. We now note that in the collinear limit the momentum in the intermediate propagator goes on shell. Indeed, we have
@@
where it is understood that the argument $p_{1}-k$ of the spinor $u_{s}$ is an on-shell momentum, $(p_{1}-k)^{2}=0$, and the denominator is expanded according to Eq. (10) about the collinear limit, in which it vanishes.
@@
Now, it should be clear that nowhere in the above argument the explicit form of $M_{0}$ was used. So what we have actually proven is that radiation of a gluon from an external quark line produces a logarithmically enhanced contribution to the square amplitude, which is proportional to the original amplitude, but with an incoming quark momentum rescaled by a factor $z\leq 1$, and a universal splitting function, which only depends on the radiation process. A similar argument also works for gluon radiation from an incoming gluon line, because one may likewise rewrite in the collinear limit the numerator of the intermediate gluon propagator as a sum over on-shell gluon polarization vectors, up to subtleties related to the choice of gauge in the nonabelian case. Of course, in this case the splitting function will be different: $P^{r}_{gg}$ now depending on the $g(p_{1})\to g(p_{1}-k)g(k)$ matrix element. Furthermore, what is being used in the argument is the on-shell behavior of the intermediate propagator. This means that the argument also works in the off-diagonal case, except that in this case an incoming quark is replaced by a gluon with factor $P^{r}_{gq}(z)$ and an incoming gluon is replaced by a quark with factor $P^{r}_{qg}(z)$. Finally, it is clear that the argument also works for massive quarks, though in this case the logarithmic integral at the lower end is regulated by the quark mass.
@@ 3.1 The Sudakov double log
In order to understand the origin of the double soft-collinear log, consider again the full set of gluon emission corrections to the tree-level Drell-Yan process (see Fig. 3). We are specifically interested in the form of the phase space integration over the emitted gluon momentum $k$. To this purpose, it is convenient to use a slightly different version of the Sudakov parametrization, namely the light-cone parametrization, in which instead of $p_{1}$ and $p_{2}$ we adopt as basis vectors their linear combinations
@@
Because we are interested in the limit in which $k$ is both collinear and soft, we can use the eikonal approximation Eq. (17). We include all first-order real emission corrections, i.e. those corresponding both to radiation from the quark and the antiquark, because the inclusion of the full gauge-invariant set of diagrams avoids some subtleties that would instead arise when computing directly the splitting function corresponding to the emission from a single quark or antiquark line, as in Sect. 2.1-2.2. Of course, the full set of diagrams also involves virtual loop corrections that to this order interfere with the leading-order process: we will come to these in Sect. 3.3.
@@
$M_{1}=M_{1}^{q}+M_{1}^{\bar{q}}=M_{0}t^{a}g_{s}\left(\frac{p_{2}^{\mu}}{p_{2}\cdot k}-\frac{p_{1}^{\mu}}{p_{1}\cdot k}\right)\epsilon^{*}_{\mu}(k)\left[1+O(1-x)\right],$ (25)
@@
$|M_{1}|^{2}=|M_{0}|^{2}C_{F}g^{2}_{s}\frac{2p_{1}\cdot p_{2}}{(p_{1}\cdot k)\,(p_{2}\cdot k)}\left[1+O(1-x)\right],$ (26)
@@
Using Eq. (32) it is apparent that only the second term in Eq. (30) contributes to the integral, which is then straightforwardly giving a squared logarithm (usually called double log). The upper limit of integration is the maximum allowed value of $k_{t}^{2}$, which is proportional to the center-of-mass energy $s$, while at the lower limit the integral diverges, so we must introduce some infrared cutoff $\mu^{2}$ in order to make sense of it. The treatment of this double divergence – one collinear and one soft – will be the subject of the next two sections. We thus get
$\int d\Phi_{k}|M_{1}|^{2}=C_{F}|M_{0}|^{2}\frac{\alpha}{2\pi}\ln^{2}\frac{s}{\mu^{2}}+\hbox{less singular}\,$, (33)
where we have neglected all contributions that are less singular as $\mu^{2}\to 0$. This includes both the O(1x)O(1-x) corrections to the eikonal approximation, which would in fact be suppressed by inverse powers of the dimensionful variable $s$, which contains the large scale of the process, but also lower powers of log, which arise when expressing the maximum value of $k_{t}^{2}$ in terms of $s$. For example, for the simple $2\to 2$ process of Fig. 1 one has ${k_{t}^{2}}_{\rm max}=s/4$ so $\ln^{2}{k_{t}^{2}}_{\rm max}=\ln^{2}s-\ln 4\ln s+\text{non log}$.
@@ 3.2 Mass singularities and their factorization
The same conclusion holds if the radiating parton is off-shell, because then the denominator of the intermediate propagator, Eq. (10), contains a contribution equal to the off-shellness $p_{1}^{2}$, and thus it does not vanish when $k_{t}^{2}\to 0$.
@@ 3.3 Cancellation of infrared singularities
After subtraction of the collinear singularity – which has been factorized into the PDF, where it is then regulated by the nonperturbative physics which determines the PDF – we are still left with an infrared singularity. This can be seen clearly by looking at the explicit form of the splitting function $P^{r}_{qq}$:
@@
A classic result of quantum electrodynamics, the Kinoshita-Lee-Nauenberg (KLN) theorem, guarantees that once one consistently includes loop corrections along with real emission contributions] this singularity cancels. The argument generalizes in QCD for colorless final states [17]. The cancellation is illustrated graphically in Fig. 4. The argument of Sect. 3.1 shows that the infrared divergent contribution to the cross-section can be obtained from the interference of the contributions to the amplitude coming from gluons emitted from different quark lines lines. Indeed, the infrared singularity comes from the $x\to 1$ limit of the denominator in the amplitude Eq. (28), which in turn arises from the interference of the two eikonal factors proportional to $p_{1}^{\mu}$ and $p_{2}^{\mu}$ in Eq. (25). The KLN theorem ensures that this singularity cancels against an equal and opposite singularity, which is present in the virtual contribution that corresponds to the gluon loop obtained by connecting the endpoints of the two interfering emitted real gluons. This loop correction has of course the same final state particles as the amplitude without either of the two interfering final-state real emitted gluons, and thus it interferes with it when calculating the square modulus of the amplitude, thereby producing a contribution of the same order of the square of either of the real emission terms (see Figure 4).
@@
Hence, the partonic cross-section contains a double-logarithmic contribution, coming from interplay between the leftover of the infrared singularity, that produces a $\frac{1}{1-z}$ behavior, and the leftover of the collinear singularity, that produces a $\ln{k_{t}^{2}}_{\rm max}$ behavior, which in turns is sensitive to the infrared scale $Q^{2}(1-z)^{2}$.
@@ 4.1 Momentum conservation and phase space factorization
In a multiple emission process an incoming parton with momentum $p_{1}$ emits a first gluon with momentum $k_{1}$ and the outgoing parton after the first emission has momentum given by $p_{1}-k$. This is then the incoming parton for the second emission, and so forth. Using the Sudakov parametrization Eq. (2) it is clear that after the first emission of a gluon with momentum
@@ 5 Renormalization group basics
The basic idea of renormalization is that in a quantum field theory all couplings and fields must be defined in terms of a reference scale, in such a way that all quantum fluctuations at distances shorter than that of the reference scale are absorbed in the definition of the couplings and fields. The fields and couplings must consequently be defined by relating them to a reference scale. The quantum field theory then provides relations between observable quantities, with the couplings and fields used as a way to relate observables to each other. Infinities only appear if one tries to express observable quantities in terms of unobservable parameters defined at infinitely short distance scales.
For example, one may define the electron charge in quantum electrodynamics in terms of the Coulomb force at a large distance, i.e. at the scale of the (rest) electron mass. The (say) elastic electron-electron scattering cross-section is finite once re-expressed in terms of the electron charge thus defined. Renormalization group invariance is then the consequence of the fact that physical predictions do not depend on the renormalization scale that has been chosen to define the couplings. Because the dependence of the couplings on this scale is a universal property of the theory, this independence expresses the scale dependence of physical observables in terms of this universal scale dependence of the coupling.
The basic idea of deriving collinear exponentiation from renormalization group invariance, that we will present shortly, takes this one step further, based on the observation that in QCD physically observable cross-sections are expressed in factorized form in terms of parton distributions, convoluted with partonic cross-sections. This factorization requires introducing a scale, as discussed in Sect. 3.2 for the subtraction of collinear singularities. As discussed in Sect. 4.2, this can be viewed as a scale that splits in two the parton radiation process, with radiation below the scale included in the parton distribution, and radiation above it included in the hard matrix element. Physical predictions are independent of this scale, hence this scale dependence must cancel between the PDF and the hard cross-section: indeed we have seen that it is expressed in terms of a universal, process-independent splitting function, or rather, its Mellin transform, the anomalous dimension Eq. (45). But the PDF and the hard cross-section depend on different variables: the hard cross-section depends on the specific hard scale of the process, while the PDF does not. As we shall see, dimensional analysis then implies that the cancellation of the dependence on the factorization scale can only go through the scale dependence of the coupling constant: it is calculable in terms of it, and the universal anomalous dimension. It can then be integrated away, leading to all-order resummation and exponentiation of logs of the hard physical scale.